OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line which passes through the points (1, 6) and (2, 9).

OpenStudy (anonymous):

\[slope=\frac{\Delta y}{\Delta x}\] Equation of a line: \[y-y_1=m(x-x_1)\], which you'll have to rearrange so that it is in the form \[y=mx+c\]

OpenStudy (anonymous):

\[9-6/2-1 \] so you get 3/1or 3 for the slope or "M"

OpenStudy (anonymous):

then you get y-6=3(x-1) OR

OpenStudy (anonymous):

y-9=3(x-2)